# Charm associated production of the pseudoscalar Higgs boson $h^{\prime\prime}$ in the MCPM’ at the LHC

### v1

@@ Appendix B Digression on total cross sections
This follows the recommendations of Ellis:1996mzs. The evolution of the proton wave function can also be considered as follows in a simple physical picture; see Kogut:1974ni; Kogut:1973ub and also sections 19.2 and 19.3 of Nachtmann:1990ta. As we increase the energy of the $pp$ collision we increase the temporal resolution. Fluctuations with higher and higher transverse momenta have to be included in the effective proton wave functions. Thus, in Fig. 20 $c$ and $\bar{c}$ quarks with transverse momenta
@@
Below we compare the obtained integrated cross sections of (10)–(14) for associated $h^{\prime\prime}+c\bar{c}$ production to the corresponding total cross sections for $h^{\prime\prime}$ production (see e.g. numbers given in (37) and (38) of Maniatis:2023aww for $\sqrt{s}=13$ TeV). For the current LHC energy $\sqrt{s}=13.6$ TeV and with the MSHT20lo_\!\_\!as130 PDFs, we find
@@
for the Drell-Yan process and $GG$ fusion, respectively. These numbers are obtained in the full phase space, i.e., without imposing any cuts on $h^{\prime\prime}$ transverse momentum. The cross sections (35) and (36) differ somewhat from those given in (37) and (38) of Maniatis:2023aww for $\sqrt{s}=13$ TeV:
@@
The uncertainty of this number is estimated to be around 20%. With the branching ratio $\text{B}(h^{\prime\prime}\to\gamma\gamma)$ as calculated in Eq.(28) of Maniatis:2023aww,
@@ Appendix C Double $h^{\prime\prime}$ production
The $h^{\prime\prime}$ decays to nearly 100% to $c\bar{c}$ for a mass value $m_{h^{\prime\prime}}=95.4$ GeV; see Fig. 8 of Maniatis:2009vp. Therefore, the reaction (C.1) will result in the final state $c\bar{c}c\bar{c}X$ and contribute to the general final states in (9) and (19).
In our calculation for (C.1), we shall restrict ourselves to the diagrams of Fig. 22(a) with charm-quark fusion ($q=c$), since the coupling of $h^{\prime\prime}$ to $c$ quarks is much bigger than to $s$ quarks; see (A.81) of Maniatis:2023aww. We calculate first the reaction
@@
Now we use this result as input for a standard Drell–Yan type calculation of the diagram of Fig. 22(a); see, for instance, chapter 18.5 of Nachtmann:1990ta. The result reads as follows:

### v2

@@ Appendix B Digression on total cross sections and comparison of $h^{\prime\prime}$ and $Z$ distributions
This follows the recommendations of Ellis et al. (1996). The evolution of the proton wave function can also be considered as follows in a simple physical picture; see Kogut and Susskind (1974b, a) and also sections 19.2 and 19.3 of Nachtmann (1990). As we increase the energy of the $pp$ collision we increase the temporal resolution. Fluctuations with higher and higher transverse momenta have to be included in the effective proton wave functions. Thus, in Fig. 20 $c$ and $\bar{c}$ quarks with transverse momenta
@@
Below we compare the obtained integrated cross sections of (10)–(14) for associated $h^{\prime\prime}+c\bar{c}$ production to the corresponding total cross sections for $h^{\prime\prime}$ production (see e.g. numbers given in (37) and (38) of Maniatis and Nachtmann (2025) for $\sqrt{s}=13$ TeV). For the current LHC energy $\sqrt{s}=13.6$ TeV and with the MSHT20lo_\!\_\!as130 PDFs, we find
@@
for the Drell-Yan process and $GG$ fusion, respectively. These numbers are obtained in the full phase space, i.e., without imposing any cuts on $h^{\prime\prime}$ transverse momentum. The cross sections (35) and (36) differ somewhat from those given in (37) and (38) of Maniatis and Nachtmann (2025) for $\sqrt{s}=13$ TeV:
@@
The uncertainty of this number is estimated to be around 20%. With the branching ratio $\text{B}(h^{\prime\prime}\to\gamma\gamma)$ as calculated in Eq.(28) of Maniatis and Nachtmann (2025),
@@
We find it interesting to compare the DY cross section (35) to the cross section for the production of the $Z$ boson with its subsequent decay to $c+\bar{c}$:
$p+p\to(Z\to c+\bar{c})+X.$ (42)
From Navas and others (2024) we have
$m_{Z}$ $=91.1880\pm 0.0020\;\text{GeV,}$
$\Gamma_{Z}$ $=2.4955\pm 0.0023\;\text{GeV,}$
$B(Z\to c+\bar{c})$ $=(12.03\pm 0.21)\times 10^{-2}.$ (43)
The mass $m_{Z}$ is close to the $m_{h^{\prime\prime}}$ mass (5). With the same methods as we calculated (35) we find
$\left.\sigma_{\mathrm{tot}}(p+p\to Z+X)\right|_{\text{DY}}=45752.7\penalty 10000\ \text{pb.}$ (44)
$\left.\sigma_{\mathrm{tot}}(p+p\to Z+X)\right|_{\text{DY}}\times B(Z\to c+\bar{c})=5490.3\penalty 10000\ \text{pb.}$ (45)
Thus, the cross section for the reaction (42) is much smaller than that of (35). Note that $h^{\prime\prime}$ boson of mass (5) is predicted to decay to nearly 100% to $c\bar{c}$; see Fig. 8 of Maniatis and Nachtmann (2009). Here we show the distributions in the longitudinal momentum $k_{3}$ and rapidity $y$ of the $h^{\prime\prime}$ in (35) and the $Z$ in (42); see Figs. 22(a) and 22(b), respectively .
Figure 22: Distribution of: (a) the longitudinal momentum $k_{3}$, and (b) the rapidity $y$ of $h^{\prime\prime}$ in (35) and of $Z$ in (42).
These distributions are not directly observable but have to be reconstructed from the distribution of the $c\bar{c}$ pairs. The invariant mass of the pairs $M_{c\bar{c}}$ is a superposition of two Breit–Wigner distributions around the nominal values of the masses of $h^{\prime\prime}$ and $Z$.
Figure 23: The distributions of $M_{c\bar{c}}$ for $h^{\prime\prime}$ production in Eq. (35) and for $Z$ production in Eq. (42), obtained (a) in the full phase space and (b) with the requirement $|y|<2.5$.
In Fig. 23 we show the distribution of $M_{c\bar{c}}$. Two peaks for the $h^{\prime\prime}$ and the $Z$ are clearly visible. Fig. 23(a) corresponds to the calculations for the full phase-space, while Fig. 23(b) represents results obtained within the $|y|<2.5$ requirement (which is equivalent to $|k_{3}|\lesssim 650$ GeV). The corresponding QCD $c\bar{c}$-background in the invariant-mass region around the two signal peaks is approximately two orders of magnitude larger and therefore, it is not shown in the figure. Note that all these distributions are calculated at the parton level. In the real world the momenta of $c$ and $\bar{c}$ have to be obtained from the corresponding jet momenta. This certainly will introduce a smearing of the distributions.
Finally we compare the distributions for the reactions
$p+p\;\to\;(h^{\prime\prime}\to\mu^{+}\mu^{-})+X,$ (B.17)
and
$p+p\;\to\;(Z\to\mu^{+}\mu^{-})+X.$ (B.18)
From Eq. (3.2) and Table 2 of Maniatis and Nachtmann (2009) we get for the $h^{\prime\prime}$ mass ($1.5$) the branching ratio of $h^{\prime\prime}\to\mu^{+}\mu^{-}$ as
$B(h^{\prime\prime}\to\mu^{+}\mu^{-})=3.44\times 10^{-5}.$ (B.19)
The branching ratio $Z\to\mu^{+}\mu^{-}$ is according to Navas and others (2024)
$B(Z\to\mu^{+}\mu^{-})=(3.3662\pm 0.0066)\times 10^{-2}.$ (B.20)
In Fig. 24, we present the analogue of Fig. 22 for the reactions (B.17) and (B.18). We observe that, in the $\mu^{+}\mu^{-}$ channel, the contribution from $Z$ production exceeds that from $h^{\prime\prime}$ by approximately three orders of magnitude, reflecting the relative difference between the corresponding branching fractions given in (B.19) and (B.20). Consequently, in this channel the invariant-mass spectrum is entirely dominated by $Z$-boson production (see Fig. 25).
Figure 24: Distribution of: (a) the longitudinal momentum $k_{3}$, and (b) the rapidity $y$ of $h^{\prime\prime}$ in (B.17) and of $Z$ in (B.18).
Figure 25: The distributions of $M_{\mu^{+}\mu^{-}}$ for Z production in Eq. (B.18) obtained in the full phase space.
In Fig. 26 we compare associated $h^{\prime\prime}$ production (6) and the corresponding associated $Z$ production
$p+p\;\to\;(Z\to c+\bar{c})+c+\bar{c}+X.$ (B.21)
Figure 26: The same as in Fig. 7(b) but here extra contribution from the reaction (B.21) is also shown (green solid histogram; color online).
We see that with the cuts (15) the contribution of (B.21) is much smaller than that of our signal process (6).
@@ Appendix C Double $h^{\prime\prime}$ production
The $h^{\prime\prime}$ decays to nearly 100% to $c\bar{c}$ for a mass value $m_{h^{\prime\prime}}=95.4$ GeV; see Fig. 8 of Maniatis and Nachtmann (2009). Therefore, the reaction (C.1) will result in the final state $c\bar{c}c\bar{c}X$ and contribute to the general final states in (9) and (19).
In our calculation for (C.1), we shall restrict ourselves to the diagrams of Fig. 27(a) with charm-quark fusion ($q=c$), since the coupling of $h^{\prime\prime}$ to $c$ quarks is much bigger than to $s$ quarks; see (A.81) of Maniatis and Nachtmann (2025). We calculate first the reaction
@@
Now we use this result as input for a standard Drell–Yan type calculation of the diagram of Fig. 27(a); see, for instance, chapter 18.5 of Nachtmann (1990). The result reads as follows: